Solve the linear system: 2x + 7y = 41, 2y = 3x + 1
Answer by checkley77(12844)

2x+7y=41 multiply this equation by 3
2y=3x+1
-3x+2y=1 multiply this equation by 2 and add them
6x+21y=123
-6x+4y=2
-----------------------
25y=125
y=125/5
y=5 ans.
2x+7*5=41
2x+35=41
2x=41-35
2x=6
x=6/2
x=3 ans.
Proof:
-3*3+2*5=1
-9+10=1
1=1